[Hallc_running] [New Logentry] RC Daily Update

simona at jlab.org simona at jlab.org
Tue Oct 9 14:25:03 EDT 2018

Logentry Text:
We took data until this morning at 7 am when Ops took beam away for RF recovery. We got to 51% of the desired statistical goal at Q2=3. W=3.14 x=0.25 theta_Shms=12.42 by that point. We will stay at this kinematics to take more data to reach 65 % of the desired stat. goal. Then we will change SHMS angle to 9.42 deg and will continue taking data.

Hall C went to controlled access this morning so that the HMS dipole valve JT4 that controls the LN2 level can be fixed. The valve is working now: https://logbooks.jlab.org/entry/3606407

There was also a walkthrough completed for 2 UT students.  

A camera was added to check the 3H04 BCM temperature: https://logbooks.jlab.org/entry/3606504

Few from Ops replaced dump diffuser cards.

Hall C requested to go to beam permit around 1 pm. We are now waiting for beam.

Other HCLOG entries that may be of interest: https://logbooks.jlab.org/entry/3606535, https://logbooks.jlab.org/entry/3606512

For a status of the data taken so far check https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page

We will meet today at 4:30 pm on the second floor of the counting house.

 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3606557
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181009/94e7cc2d/attachment.html>

More information about the Hallc_running mailing list